How can i get the numbers 20,17,39,and 13 by using 5,2,8,and 4 only once.I can only add,subtract,and multiply once.
Answer by jim_thompson5910(28717)

Generating 20:

(5+2)*4-8 = 7*4-8 = 28-8 = 20

or

(5-2)*4+8 = 3*4+8 = 12+8 = 20

or

(8-5+2)*4 = 5*4 = 20



Generating 17:
5+2*8-4 = 5+16-4 = 21-4 = 17


Generating 39:
It is NOT possible to generate 39 using the given numbers and rules above. I've tried all possible combinations and cannot get 39 (you could get 39 if you allowed two + symbols since 8*4+5+2 = 39)


Generating 13
(8-4)*2+5 = 4*2+5 = 8+5 = 13